OpenStudy (anonymous):

How do I solve arctan[cos(7pi/2)]? Cos of 7pi/2 isn't on the unit circle.

OpenStudy (anonymous):

you know 7pi/2 =3pi/2 right

OpenStudy (anonymous):

What? 6 pi/2 = 3/2... But not 7 pi... Right..? I don't understand.

OpenStudy (anonymous):

you know that cos(x) = cos(x + k(2pi)) right?

OpenStudy (anonymous):

what do you think cos(7pi/2) equals to in terms of cos( *number here* ) ?

OpenStudy (anonymous):

Should I just add 2pi? 21 pi / 2..?

OpenStudy (anonymous):

well if you add, though the value of cosine stays the same, the angle is bigger. Do you think that would help? what is the range of angle we usually use the most?

OpenStudy (anonymous):

0-2pi

OpenStudy (anonymous):

is 7pi/2 in that range?

OpenStudy (anonymous):

I still don't understand where to plug in 21 pi / 2 though

OpenStudy (anonymous):

Yeah it is

OpenStudy (anonymous):

Oh wait, I need to subtract 2 pi?

OpenStudy (anonymous):

BINGO!

OpenStudy (anonymous):

ohhhh. 3pi/2

OpenStudy (anonymous):

That makes a lot more sense. Thank you

OpenStudy (anonymous):

so cos(7pi/2) = cos(7pi/2 - 2pi) yes?

OpenStudy (anonymous):

Yeah. So that means after I find the cosine value and substitute it for the inverse tangent, the answer radian is undefined?

OpenStudy (anonymous):

nope. What did you get for cos(3pi/2) ?

OpenStudy (anonymous):

Oh, 0.

OpenStudy (anonymous):

excellent. What is arctan(0) ?

OpenStudy (anonymous):

0+pi k?

OpenStudy (anonymous):

well no, 0 in this case is no longer the angle, why?

OpenStudy (anonymous):

it's the radian measure?

OpenStudy (anonymous):

Is it just 0?

OpenStudy (anonymous):

correct it's just 0 (a number). For example, sin(pi/6) = 1/2. pi/6 is the angle. What is 1/2 then? well, just a number for sine when the angle is pi/6. Make sense?

OpenStudy (anonymous):

Yeah, so I wouldn't have to call it a radian or an angle measure... But isn't Pi another number that answers arctan (0)? 0/-1 = 0...

OpenStudy (anonymous):

you're right. As soon as we evaluate arctan(0), which equals 0, <-- this 0 is actually now the angle in radian. just like, arctan(1/sqrt(2)) = pi/4, <-- this is the angle. arctan(0) = 0 < this is the angle ..........^ this is not.

OpenStudy (anonymous):

OK that makes sense. But what if I was asked to find the angles where tan = a value, like 3.7321?

OpenStudy (anonymous):

Is that the same sort of problem, solving for angles?

OpenStudy (anonymous):

Unfortunately, finding the angle by hand is subject to limitation. There are special cases like cos(Θ) = 1/2, tan(Θ) = 1/sqrt(2) ... where we can easily find the angles using the unit circle. However, given any random number k, e.g. cos(Θ) = k, tan(Θ) = k, etc...it's likely that we'd have to rely on technology, like a calculator.

OpenStudy (anonymous):

Okay. What about cos x = -0.8?

OpenStudy (anonymous):

you'd need a calculator for that one

OpenStudy (anonymous):

Alright. How would I know when to add +2pi k or +pi k? For a question like cos x = sqrt(3)/2, I could find that pi/6 is the radian value, but would I add 2pi k to that problem, or just leave it at pi/6?

OpenStudy (anonymous):

well, cos(x), sin(x), tan(x) are period, so you'll need to add the period. In the case of cos(x) = sqrt(3)/2 x = pi/6 + k(2pi). There are also x values but I don't say it here

OpenStudy (anonymous):

...are *periodic*

OpenStudy (anonymous):

you know how cos(x) = cos(x + k(2pi)) ? then if x=t is the root, the ALL roots will be x = t + k(2pi)

OpenStudy (anonymous):

Okay. I get that. How can I solve for x if I've got Cos^2 (x) = 1/2?

OpenStudy (anonymous):

how would you get rid of the square?

OpenStudy (anonymous):

Take the square root. So Sqrt 2 / 2?

OpenStudy (anonymous):

close. What happens when you take a square root?

OpenStudy (anonymous):

+ or -?

OpenStudy (anonymous):

very good. So what do you have?

OpenStudy (anonymous):

+ or - sqrt 2 / 2....+ 2 pi k?

OpenStudy (anonymous):

no, sqrt(2)/2 is just the number. You only add k(2pi) to the roots.

OpenStudy (anonymous):

oh okay. What about a more complicated equation, like sec^2 (x) + 3 tan^2 (x) = 13 I can already tell that I want to use the Pythagorean identity that Tan^2 (x) + 1 = Sec^2 (x), but I can't tell where to start the problem. Dividing out the 3 would just create problems with the Secant, wouldn't it?

OpenStudy (anonymous):

The truth is that you can't possibility solve every single trig question out there. There are many messy trig equations that, when trig identities are used, amazingly reduced to the ones we can work with. But that is still limited. So the case of the function above, you well need to use a graphing calculator.

OpenStudy (anonymous):

I just realized we can solve the equation above by hand O.O. But then still, say you have something like cos^3(x) - sin(5x) cot^3(x) - 1 = 0, then a graphing calculator is guaranteed.

OpenStudy (anonymous):

I see. But on this problem, where would I start? Dividing the 3 out would be wrong, would it not?

OpenStudy (anonymous):

this is an example where a trig identity is used to reduced a messy equation to a nice equation. sec^2(x) = tan^2(x) + 1 sec^2 (x) + 3 tan^2 (x) = 13 [tan^2(x) + 1] + 3tan^2(x) = 13 4tan^2(x) + 1 = 13 4tan^2(x) = 12 tan^2(x) = 3 as you can see, that's pretty nice.

OpenStudy (anonymous):

Now I see it, is there any thing else I can do afterwards to solve for x in the final equation? Tan^2(x)=3?

OpenStudy (anonymous):

take the square root

OpenStudy (anonymous):

tan x = sqrt (3) What now?

OpenStudy (anonymous):

tan(x) = sqrt(3) or tan(x) = -sqrt(3)

OpenStudy (anonymous):

can I find x?

OpenStudy (anonymous):

tan(x) = sqrt(3) what angle would that be?

OpenStudy (anonymous):

pi/3?

OpenStudy (anonymous):

there is one more within [0,2pi]

OpenStudy (anonymous):

4pi/3?

OpenStudy (anonymous):

good. What is the period of tan(x) ?

OpenStudy (anonymous):

pi k so, +pi k

OpenStudy (anonymous):

One more question, I need to simplify this equation. I'm pretty sure it simplifies by hand. Sin^2 (x) + sin x - 2 = 0

OpenStudy (anonymous):

correct, x = pi/3 + k(pi) or x = 4pi/3 + k(pi) you exactly the same in the case of tan(x) = -sqrt(3)

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

can you factor u^2 + u - 2 ?

OpenStudy (anonymous):

OH it's factoring I didn't even catch that

OpenStudy (anonymous):

yes, this is why it's really also important to know how factor when you encounter a problem like this

OpenStudy (anonymous):

I would have to use Quadratic formula, right?

OpenStudy (anonymous):

You could, though it can be factored as (u - 1) (u+2)

OpenStudy (anonymous):

are you sure? Isn't there a coeffecient of two in front of u^2 though? I don't follow...

OpenStudy (anonymous):

I believe you're were hallucinating :P